# ToTRL: Unlock LLM Tree-of-Thoughts Reasoning Potential through Puzzles Solving

### v1

@@ 3.2 In-Distribution Logic Reasoning
3.3 Out-of-Distribution Logic Reasoning
Besides in-distribution tasks, we also evaluate our ToTQwen3-8B model on OOD logic reasoning tasks. As shown in Table 2 and Table 3, many models exhibit very low scores (close to 0), indicating a significant struggle with these logic reasoning tasks.
Logic Reasoning with Unique Solution. Table 2 presents performance on OOD logic reasoning tasks, all of which have a single correct solution. It is important to note that for tasks with unique solutions, some models might occasionally arrive at the correct answer through guessing rather than a complete logical derivation process. Our ToTQwen3-8B consistently achieves the highest scores across all three unique solution tasks evaluated, 0.508 on 5$\times$5 Crossword, 0.260 on 9$\times$9 Sudoku, and 0.986 on K&K Puzzle. This consistent, high performance on diverse OOD tasks provides strong evidence that our ToTQwen3-8B model effectively leverages the underlying logical constraints and employs a well-developed derivation process, rather than relying on chance.
Logic Reasoning with Multiple Solutions. Table 3 presents performance on more OOD logic reasoning tasks, including Poker 24 Game and Make 24 Puzzle, which admit multiple valid solutions. The primary challenge lies not only in identifying a valid solution but also potentially in finding multiple distinct solutions or performing robustly across puzzles with varying numbers of possible solutions. This necessitates a reasoning process capable of exploring diverse successful paths rather than merely converging on a single outcome. As shown in Table 3, ToTQwen3-8B again outperforms all other models on both the Poker 24 Game and Make 24 Puzzle, demonstrating superior performance across puzzles with varying numbers of solutions. This advantage is most pronounced on puzzles with many solutions, where the capacity to explore multiple valid reasoning paths is essential. The ToT strategy inherently explores a tree of possibilities, branching at critical decision points. This systematic exploration of multiple potential continuations provides a global perspective, making it inherently well-suited for tasks admitting diverse valid solutions. Unlike long CoT reasoning strategy that commits to a single path, ToT can explore parallel branches, significantly increasing the probability of discovering multiple distinct solutions when they exist.
In summary, from an OOD perspective, the results presented in Table 2 and Table 3 strongly indicate that ToTQwen3-8B possesses superior generalization capabilities on these logic reasoning tasks. Its robust performance on both unique and multiple-solution puzzles suggests that the ToT reasoning structure confers a significant advantage in addressing unfamiliar logic challenges by facilitating a more comprehensive exploration of reasoning possibilities, a capability that is particularly valuable for tasks requiring the discovery of multiple valid solutions.
@@ 3.3 Out-of-Distribution Logic Reasoning
3.4 Out-of-Distribution Mathematical Tasks
We further investigate whether our ToTRL can enhance complex reasoning abilities can transfer to a highly challenging mathematical reasoning scenario. According to Table 4, ToTQwen3-8B demonstrates strong performance on these OOD mathematical reasoning tasks. It achieves the highest average score of 0.750 across the three benchmarks. Specifically, ToTQwen3-8B performs comparably well on AIME 2024 with a score of 0.667, matching the performance of GLM-4-Z1-9B-0414 [28] and Phi-4 Reasoning [1] with more model parameters. On AIME 2025, ToTQwen3-8B leads with a score of 0.633, outperforming all other listed models. For the AMC 2023 benchmark, ToTQwen3-8B achieves a score of 0.950, which is tied with GLM-4-Z1-9B-0414 [1] for the highest performance.
ToTQwen3-8B exhibits a superior ability to handle the complex and varied mathematical problems in the AIME and AMC competitions. This strong performance on OOD tasks indicates that our ToTRL framework can effectively enhance the reasoning capabilities of LLMs. Importantly, these improved abilities generalize well to novel and challenging mathematical problems beyond the training distribution. This suggests that ToTRL successfully cultivates a more robust and transferable form of reasoning.
3.5 Test Time Scaling
Test-time scaling [25] is an innovative approach in language modeling that leverages additional computational resources during the testing phase to enhance performance. This method has shown significant promise in various domains, including language modeling and code generation [13, 9, 6, 8]. We investigate the test time scaling of our proposed ToTQwen3-8B model against a baseline Qwen3-8B [14] model using different thinking budgets across six logic reasoning tasks.
As illustrated in Figure 3, the performance of both ToTQwen3-8B and Qwen3-8B generally improves with an increased thinking budget across all tasks. This indicates that allowing more computational resources for intermediate thinking steps leads to better reasoning outcomes. Importantly, Figure 3 demonstrates the efficiency of the ToTQwen3-8B approach, which leverages the ToT reasoning strategy. ToTQwen3-8B consistently outperforms Qwen3-8B across various thinking budgets. Notably, ToTQwen3-8B is often able to reach a higher average score with a smaller thinking budget compared to Qwen3-8B. This suggests that the structured tree-of-thoughts reasoning employed by ToTQwen3-8B allows it to explore the solution space more effectively and efficiently, requiring fewer tokens (and thus less computational cost and time) to achieve superior or comparable performance. This efficiency is a key advantage, making ToTQwen3-8B a more practical solution for logic reasoning tasks under computational constraints.
Figure 3: Illustration of test time scaling on logic reasoning tasks.
@@ 3.6 Ablation Studies
To further investigate the contribution of ToT guidance prompts and the multi-stage training process within ToTRL, we perform a series of ablation studies.
ToT Guidance. Table 5 presents an ablation study on the effectiveness of ToT guidance compared to CoT guidance. For the Qwen3-8B model, employing ToT guidance yields mixed results. The performance improves on some tasks but decreases significantly on others. This suggests that although ToT guidance holds potential, the Qwen3-8B model is not adept at utilizing the ToT reasoning strategy. In contrast, when applying ToT guidance to the ToTQwen3-8B model, we observe consistent and significant improvement across all reported tasks for both guidance types. This clearly demonstrates that ToT guidance is substantially more effective than CoT, particularly when paired with a model specifically trained to leverage its tree-like exploration capabilities. Notably, ToTQwen3-8B also demonstrates improved performance with CoT guidance, indicating that ToTRL can also facilitate CoT reasoning capabilities.
Multi-Stage ToTRL. Table 6 presents an investigation into the contribution of the multi-stage training process within ToTRL, comparing models trained with different stages. The results demonstrate that the inclusion of Stage 1 training significantly improves performance across all evaluated tasks. Specifically, Stage 1 training focuses on enabling the LLM to perform the fundamental steps of ToT reasoning using no-thinking mode, thereby facilitating the integration of this capability into models primarily accustomed to sequential CoT reasoning. This indicates that the initial exploration and tree-building capabilities acquired during Stage 1 are crucial for effective ToT guidance and the superior performance on diverse reasoning tasks.
@@
CoT ToT Crossword 5$\times$5 Sudoku 9$\times$9 K&K Puzzle Poker 24 Game Make 24 Puzzle
Qwen3-8B (Enable-Thinking) 0.378 0.180 0.950 0.468 0.614
@@
Stage 1 Stage 2 Crossword 5$\times$5 Sudoku 9$\times$9 K&K Puzzle Poker 24 Game Make 24 Puzzle
@@ 4 Conclusion
In this work, we introduced ToTRL to guide LLMs from sequential CoT reasoning to a more efficient ToT reasoning strategy. By employing a two-stage training process and utilizing puzzle games that necessitate tree-based exploration, we successfully cultivate ToT capabilities in LLMs. Our ToTQwen3-8B, trained with ToTRL, demonstrates substantially superior performance on in-domain logic puzzles and generalization to OOD tasks. Furthermore, these enhanced reasoning abilities transfer effectively to challenging mathematical benchmarks. Crucially, our model also exhibits greater efficiency, achieving higher scores with smaller thinking budgets during test-time scaling compared to its CoT-based counterpart. These results collectively overcome the verbosity and local perspective limitations of long CoT, showcasing the robust potential and practical advantages of explicit tree-based reasoning for advanced AI on diverse complex tasks.
@@ A.1 Tree-of-Thoughts
Early applications of ToT [22, 11] approaches relied on external search algorithms (e.g., BFS, DFS) or auxiliary modules (e.g., prompter agents, checkers, memory modules, and ToT controllers) to manage planning, decision-making, and backtracking. Inspired by AlphaZero, TS-LLM [7] proposed learning a dedicated value function to guide the tree search and iteratively improving the LLM itself, aiming to handle deeper and more complex search trees. Collectively, these efforts demonstrate the significant potential of internalizing ToT capabilities within the LLM itself, moving towards more autonomous reasoning. However, significant challenges remain in achieving truly autonomous planning and decision-making, as existing methods frequently depend on external search control rather than fully developing the LLM’s inherent capacity for tree-based reasoning. Consequently, building upon the long-CoT reasoning capability, we develop ToTRL to train LLMs to autonomously perform ToT reasoning.
A.2 Long-COT Activated through RL
RL [16, 17, 2, 26] has proven an effective approach for eliciting longer CoT, thereby enhancing LLMs’ reasoning capabilities. DeepSeek-R1 [9] and Kimi-K1.5 [6] have demonstrated that LLMs can acquire and extend their reasoning paths, including reflection and verification, using RL with simple rule-based rewards. Data-driven RL research has further broadened the application of this approach. For instance, Logic-RL [21] utilized logic puzzles for training, demonstrating generalization to mathematics, while SWE-RL [19] leveraged software evolution data to improve model performance on software engineering tasks and OOD reasoning. However, despite somewhat enhancing the depth of thought and problem-solving capabilities, current RL-elicited long CoT [18, 11] reasoning paradigms are inherently linear, following a single exploration path. This sequential structure is inefficient for tackling complex problems that necessitate extensive exploration and evaluation of multiple potential solutions. Lacking a global perspective and effective mechanisms for path evaluation and pruning, this approach often generates redundant outputs and incurs unnecessary computational overhead. In this paper, we develop ToTRL to guide LLM reasoning from a global perspective, which aims to improve performance while reducing token costs.
Appendix B Experiment Details
@@
6$\times$6 Sudoku Self-generated Dataset 50 -
@@
5$\times$5 Crossword ToT Dataset [22] 50 -
9$\times$9 Sudoku Kaggle Sudoku Dataset [5] 50 -
K&K Puzzle K&K Puzzles Dataset [20] 140 20
@@ B.1 Training Data
As detailed in Table 7, the training data is evenly distributed between two types of self-generated puzzles, including 720 6$\times$6 Sudoku puzzles and 720 Alphametic puzzles. The training regimen is divided into two stages, with Stage 1 comprising 1080 puzzles and Stage 2 comprising 360 puzzles. Notably, the Alphametic puzzles in the training set are specifically designed to feature between 1 and 4 unique solutions, thereby ensuring the model’s exposure to a diverse range of problem complexities.
@@ B.2 Evaluation Benchmarks
The performance of the ToTQwen3-8B is rigorously evaluated across various tasks, encompassing both in-distribution puzzles and OOD challenges that the model has not encountered previously. A comprehensive overview of these evaluation benchmarks, including their categories, sources, test set sizes, and specific sample distributions, is presented in Table 8.
@@
For in-distribution evaluation, the evaluation dataset for 50 6$\times$6 Sudoku puzzles and 200 Alphametic puzzles is self-generated. The Alphametic Puzzle test set maintains a distribution of 50 puzzles per solution count category, including 1 to 4 solutions, to robustly assess performance.
@@
The OOD evaluation dataset is derived from established datasets and self-generated puzzles. Standardized puzzles include 50 5$\times$5 Crossword puzzles from the ToT dataset [22] and 50 9$\times$9 Sudoku puzzles randomly selected from a Kaggle Dataset [5] to test generalization to more complex Sudoku formats. The K&K Puzzle benchmark, sourced from the original K&K Puzzles dataset  [20], comprises 140 test puzzles, specifically chosen to provide 20 puzzles for each of its 7 distinct solution categories. To further probe multi-solution reasoning in novel contexts, we self-generated 160 puzzles each for the Poker 24 Puzzle and Make 24 Puzzle. These datasets are carefully structured to include 40 samples for problems featuring 1, 2, 3, or 4 distinct solutions, allowing for a nuanced evaluation of the model’s ability to identify multiple valid outcomes.
Appendix C Limitations
Despite the promising results of ToTQwen3-8B, this work has several limitations. The transferability of ToTQwen3-8B to different reasoning tasks requires a thorough investigation. Furthermore, our approach implicitly induces ToT reasoning through guided prompting and a two-stage reinforcement learning strategy, leveraging training data based on long CoT outputs. This method aims to encourage a more global problem perspective and potentially reduce computational overhead compared to generating exhaustive CoT traces for every thought path. However, a limitation is that the model, trained on these CoT foundations, may still produce partially redundant outputs. This redundancy can stem from the underlying long CoT structure, even with the ToT mechanism attempting to focus exploration. This implicit induction, although effective in partially mitigating CoT habituation and demonstrating the aforementioned benefits, may exhibit different exploratory dynamics and optimality compared to explicit tree search algorithms. Explicit algorithms could potentially offer stricter pruning or different trade-offs in managing redundancy.

### v2

@@ 3.6 Ablation Studies
To further investigate the contribution of ToT guidance prompts and the multi-stage training process within ToTRL, we perform a series of ablation studies. Notably, additional ablation studies are provide in Appendix˜C.
ToT Guidance. Table˜5 presents an ablation study on the effectiveness of ToT guidance compared to CoT guidance. For the Qwen3-8B model, employing ToT guidance yields mixed results. The performance improves on some tasks but decreases significantly on others. This suggests that although ToT guidance holds potential, the Qwen3-8B model is not adept at utilizing the ToT reasoning strategy. In contrast, when applying ToT guidance to the ToTQwen3-8B model, we observe consistent and significant improvement across all reported tasks for both guidance types. This clearly demonstrates that ToT guidance is substantially more effective than CoT, particularly when paired with a model specifically trained to leverage its tree-like exploration capabilities. Notably, ToTQwen3-8B also demonstrates improved performance with CoT guidance, indicating that ToTRL can also facilitate CoT reasoning capabilities.
Multi-Stage ToTRL. Table˜6 presents an investigation into the contribution of the multi-stage training process within ToTRL, comparing models trained with different stages. The results demonstrate that the inclusion of Stage 1 training significantly improves performance across all evaluated tasks. Specifically, Stage 1 training focuses on enabling the LLM to perform the fundamental steps of ToT reasoning using no-thinking mode, thereby facilitating the integration of this capability into models primarily accustomed to sequential CoT reasoning. This indicates that the initial exploration and tree-building capabilities acquired during Stage 1 are crucial for effective ToT guidance and the superior performance on diverse reasoning tasks.
@@
CoT ToT Crossword $\times5$ Sudoku $\times9$ K&K Puzzle Poker 24 Game Make 24 Puzzle
Qwen3-8B (Thinking) 0.378 0.180 0.950 0.468 0.614
@@
Stage 1 Stage 2 Crossword $\times5$ Sudoku $\times9$ K&K Puzzle Poker 24 Game Make 24 Puzzle
@@ 4 Conclusion
In this work, we introduced ToTRL to guide LLMs from sequential CoT reasoning to a more efficient ToT reasoning strategy. By employing a two-stage training process and utilizing puzzle games that necessitate tree-based exploration, we successfully cultivate ToT capabilities in LLMs. Our ToTQwen3-8B, trained with ToTRL, demonstrates substantially superior performance on in-domain logic puzzles and generalization to OOD tasks. Furthermore, these enhanced reasoning abilities transfer effectively to challenging mathematical benchmarks. Crucially, our model also exhibits greater efficiency, achieving higher scores with smaller thinking budgets during test-time scaling compared to its CoT-based counterpart. These results collectively overcome the verbosity and local perspective limitations of long CoT, showcasing the robust potential and practical advantages of explicit tree-based reasoning on diverse complex tasks.
Limitations
Despite the promising results of ToTQwen3-8B, this work has several limitations. The transferability of ToTQwen3-8B to different reasoning tasks requires a thorough investigation. Furthermore, our approach implicitly induces ToT reasoning through guided prompting and a two-stage reinforcement learning strategy, leveraging training data based on long CoT outputs. This method aims to encourage a more global problem perspective and potentially reduce computational overhead compared to generating exhaustive CoT traces for every thought path. However, a limitation is that the model, trained on these CoT foundations, may still produce partially redundant outputs. This redundancy can stem from the underlying long CoT structure, even with the ToT mechanism attempting to focus exploration. This implicit induction, although effective in partially mitigating CoT habituation and demonstrating the benefits mentioned above, may exhibit different exploratory dynamics and optimality compared to explicit tree search algorithms. Explicit algorithms could potentially offer stricter pruning or different trade-offs in managing redundancy.
@@ A.1 Tree-of-Thoughts
Early applications of ToT (yao2023tot; long2023tot) approaches relied on external search algorithms (e.g., BFS, DFS) or auxiliary modules (e.g., prompter agents, checkers, memory modules, and ToT controllers) to manage planning, decision-making, and backtracking. Inspired by AlphaZero, TS-LLM (feng2023tsllm) proposed learning a dedicated value function to guide the tree search and iteratively improving the LLM itself, aiming to handle deeper and more complex search trees. Furthermore, ATS (zhang2023ats) allows LLMs to perform complex tree-search reasoning by generating the entire search trajectory in a single response, which can be activated by prompting LLMs. Collectively, these efforts demonstrate the significant potential of internalizing ToT capabilities within the LLM itself, moving towards more autonomous reasoning. However, significant challenges remain in achieving truly autonomous planning and decision-making. Existing methods frequently depend on external search control and prompting rather than developing the LLM’s inherent capacity for tree-based reasoning. Consequently, building upon the long-CoT reasoning capability, we develop ToTRL to train LLMs to perform ToT reasoning.
@@
$\times6$ Sudoku Self-generated Dataset 50 -
@@
$\times5$ Crossword ToT Dataset (yao2023tot) 50 -
$\times9$ Sudoku Kaggle Sudoku Dataset (kaggle9dataset) 50 -
K&K Puzzle K&K Puzzles Dataset (xie2024memorization) 140 20
@@
A.2 Long-COT Activated through RL
RL (schulman2017ppo; shao2024deepseekmath; ahmadian2024rloo) has proven an effective approach for eliciting longer CoT, thereby enhancing LLMs’ reasoning capabilities. Recent studies (guo2025deepseekr1; kimi2025k1.5) have demonstrated that LLMs can acquire and extend their reasoning paths, including reflection and verification, using RL with simple rule-based rewards. Data-driven RL research has further broadened the application of this approach. For instance, Logic-RL (xie2025logicrl) utilized logic puzzles for training, demonstrating generalization to mathematics, while SWE-RL (wei2025swerl) leveraged software evolution data to improve model performance on software engineering tasks and OOD reasoning. However, despite somewhat enhancing the depth of thought and problem-solving capabilities, current RL-elicited long CoT (wei2022cot; long2023tot) reasoning paradigms are inherently linear, following a single exploration path. This sequential structure is inefficient for tackling complex problems that necessitate extensive exploration and evaluation of multiple potential solutions. Lacking a global perspective and effective mechanisms for path evaluation and pruning, this approach often generates redundant outputs and incurs unnecessary computational overhead. In this paper, we develop ToTRL to guide LLM reasoning from a global perspective, which aims to improve performance while reducing token costs.
Appendix B Experiment Details
@@ B.1 Training Data
As detailed in Table˜7, the training data is distributed between two types of self-generated puzzles, including 720 $\times6$ Sudoku puzzles and 720 Alphametic puzzles. The training regimen is divided into two stages, with stage 1 comprising 1080 puzzles and stage 2 comprising 360 puzzles. Notably, the Alphametic puzzles in the training set are specifically designed to feature between 1 and 4 unique solutions, thereby ensuring the model’s exposure to a diverse range of problem complexities.
@@ B.2 Evaluation Benchmarks
The performance of the ToTQwen3-8B is rigorously evaluated across various tasks, encompassing both in-distribution puzzles and OOD challenges that the model has not encountered previously. A comprehensive overview of these evaluation benchmarks, including their categories, sources, test set sizes, and specific sample distributions, is presented in Table˜8.
@@
For in-distribution evaluation, the evaluation dataset for 50 $\times6$ Sudoku puzzles and 200 Alphametic puzzles is self-generated. The Alphametic Puzzle test set maintains a distribution of 50 puzzles per solution count category, including 1 to 4 solutions, to robustly assess performance.
@@
The OOD evaluation dataset is derived from established datasets and self-generated puzzles. Standardized puzzles include 50 $\times5$ Crossword puzzles from the ToT dataset (yao2023tot) and 50 $\times9$ Sudoku puzzles randomly selected from a Kaggle Dataset (kaggle9dataset) to test generalization to more complex Sudoku formats. The K&K Puzzle benchmark, sourced from the original K&K Puzzles dataset  (xie2024memorization), comprises 140 test puzzles, specifically chosen to provide 20 puzzles for each of its 7 distinct solution categories. To further probe multi-solution reasoning in novel contexts, we self-generated 160 puzzles each for the Poker 24 Puzzle and Make 24 Puzzle. These datasets are carefully structured to include 40 samples for problems featuring 1-4 distinct solutions, allowing for a nuanced evaluation of the model’s ability to identify multiple valid outcomes.
Table 9: Ablation study on ToT thinking across various LLMs.
Crossword $\times5$ Sudoku $\times9$ K&K Puzzle Poker 24 Game Make 24 Puzzle
DeepSeek-R1-Distill-Qwen-7B 0.000 0.000 0.012 0.010 0.000
Llama-3.1-Nemotron-Nano-8B 0.000 0.000 0.032 0.047 0.033
GLM-4-Z1-9B-0414 0.060 0.000 0.710 0.356 0.417
Phi-4 Reasoning 0.000 0.040 0.722 0.054 0.232
Qwen3-8B (Thinking) 0.376 0.080 0.700 0.485 0.600
ToTQwen3-8B (Ours) 0.508 0.260 0.986 0.603 0.661
Table 10: Ablation study on reward modeling.
Reward Modeling Crossword $\times5$ Sudoku $\times9$ K&K Puzzle Poker 24 Game Make 24 Puzzle
Qwen3-8B Partial-Credit 0.200 0.040 0.836 0.238 0.315
Full-Credit 0.508 0.260 0.986 0.603 0.661
Table 11: Ablation study on different training paradigms.
Training Paradigm Crossword $\times5$ Sudoku $\times9$ K&K Puzzle Poker 24 Game Make 24 Puzzle
Qwen3-8B CoT 0.404 0.180 0.957 0.526 0.628
ToT 0.508 0.260 0.986 0.603 0.661
B.3 Thinking Budget
The thinking budget for the model’s intermediate thinking output is predefined (qwen2025qwen3). Should the thinking output reach the thinking budget, the thinking process is terminated, and a standardized instruction is immediately introduced: “Considering the limited time by the user, I have to give the solution based on the thinking directly now.” The model then generates its final response based on the reasoning accumulated up to the termination point. This procedure guarantees that all methods compared operate under an equivalent and constrained computational budget.
Appendix C Supplementary Ablation Studies
Thinking with ToT. As indicated in Table˜9, applying the ToT reasoning strategy via prompting to the untrained baseline models resulted in performance that is generally inferior or highly inconsistent compared to their native CoT reasoning. This finding strongly validates the core motivation of this work: untrained LLMs cannot effectively utilize the ToT strategy through simple prompting alone. Specialized training with ToTRL is essential to unlock and maximize this potential.
Reward Modeling. Table˜10 presents an exploration of partial-credit rewards within the ToTRL framework. As shown, employing partial rewards caused the model to prematurely converge on sub-optimal strategies that only output a subset of the correct answers. This ultimately collapsed the intended ToT thinking process and yielded lower overall scores compared to the full-credit approach.
Training Paradigms. Table˜11 presents an ablation study comparing the effectiveness of ToT guidance against CoT guidance under an identical total number of training steps. As demonstrated, the ToT thinking paradigm provides a clear performance improvement over the CoT paradigm. Consequently, the observed gain cannot be attributed merely to extended training.
Table 12: Comparison to explicitly guided search algorithms on poker 24 game.
   Game    CoT    ToT    TS-LLM    ToTRL
   Poker 24 Game    0.468    0.631    0.582    0.603
Table 13: Performance of ToTQwen3-8B on OOD real-world tasks.
Game BFCL v3 (Live) LiveCodeBench Arena Hard Creative Writing v3
Qwen3-8B (Thinking) 0.767 0.556 0.832 0.727
ToTQwen3-8B (Ours) 0.783 0.554 0.858 0.756
Appendix D Case Studies
We analyze the following alphametic puzzle: $\text{QQB}+\text{QVQ}=\text{VBG}$, where each letter represents a distinct digit and leading zeros are disallowed. The objective is to enumerate all possible assignments that satisfy this equation. Figure˜4 illustrates the reasoning trajectory and final solutions produced by the Qwen3-8B using a standard CoT prompt, while Figure˜5 displays the tree-structured reasoning process of ToTQwen3-8B under a ToT prompt. We utilize this specific example to compare the behavior of these two distinct reasoning approaches on the same task.
The two approaches exhibit notable differences in their reasoning structure. As shown in Figure˜4, the CoT model develops its reasoning along a single, linear path. It documents local column-wise addition and carry constraints while repeatedly alternating between different hypotheses for the values of Q and V within that same path. Consequently, the distinct search branches become entangled, lacking clear case partitioning. This structural deficiency makes checks across branches prone to redundancy and omissions. In contrast, ToTQwen3-8B in Figure˜5 initially partitions the possible values of Q into several cases based on the constraint in the hundreds column, and then further divides these into subcases according to the value of V. Within each resulting branch, the model sequentially applies the constraints from the units, tens, and hundreds columns. As soon as a contradiction emerges in any column, that branch is immediately pruned and is not expanded further, yielding a structurally clear search tree that contains substantially less redundant exploration.
The disparity in reasoning structure is further evident in the quality of the outputs. In Figure˜4, the CoT model’s “<answer>” block provides three assignments. Although (Q, V, B, G)=(1,2,3,4)($\text${Q, V, B, G})=(1,2,3,4) and $(2,4,6,8)$ satisfy all column-wise addition and digit-uniqueness constraints, the assignment $(2,4,7,9)$ is inconsistent with the global constraints yet is incorrectly labeled as another correct solution. Furthermore, the valid solution $(4,9,3,7)$ is entirely missed. This demonstrates that, in the linear CoT mode, the model’s global verification is unreliable, especially in multi-solution settings. Conversely, in Figure˜5, the ToT model explicitly differentiates valid solutions from no-solution branches throughout the tree and ultimately aggregates the three correct assignments: $(1,2,3,4)$, $(2,4,6,8)$, and $(4,9,3,7)$. These assignments precisely constitute all true solutions to the puzzle, with no missing, incorrect, or duplicate entries.
Appendix E Discussion
E.1 ToT Activation
Similar to how long CoT reasoning process is often activated and generalized via RL training on mathematical tasks (guo2025deepseekr1), our contribution demonstrates that an appropriate task set (specifically, puzzles) can effectively activate the ToT thinking mode. As indicated in Table˜13, we observe modest performance improvements on agentic tool use (BFCL v3 (yan2024bfcl)) and writing tasks (Arena Hard (li2024arenahard), Creative Writing (creativewriting)), while no statistically significant difference has been observed so far on coding benchmarks (LiveCodeBench v5 (jain2024livecodebench)). These findings suggest that our approach holds preliminary yet promising potential for real-world applications. Future work can focus squarely on generalizing this learned exploration policy using RL to real-world scenarios.
E.2 Explicitly-Guided Search Strategies
Explicitly guided search algorithms typically necessitate customized frameworks for different tasks. In contrast, our method avoids explicitly guiding the model to use a specific search algorithm (e.g., Breadth-First Search (BFS) or Monte Carlo Tree Search (MCTS)); instead, it trains a policy that learns to explore. To demonstrate the performance disparity between these approaches, we compare ToTRL with ToT (yao2023tot) and TS-LLM (feng2023tsllm) on the Poker 24 Game, utilizing the Qwen3-8B-think model for both ToT and TS-LLM. As shown in Table˜12, our ToTRL achieves performance similar to these explicitly guided tree-search methods, but offers greater practicality. This improved practicality stems from requiring only a single invocation and eliminating the need for a task-specific framework. This result effectively demonstrates the strong utility of internalizing the ToT thinking mode into LLMs.
Figure 4: Illustration of CoT reasoning pattern with solving an Alphametic puzzle as an example.
Figure 5: Illustration of ToT reasoning pattern with solving an Alphametic puzzle as an example.